x^2 = 16



y^2 + 8y = 0

Answer by stanbon(75887)

x^2 = 16
x^2-16 = 0
Factor:
(x-4)(x+4) = 0
x = 4 or x = -4
=========================

y^2 + 8y = 0
Factor:
y(y+8) = 0
y = 0 or y = -8
=====================
Cheers,
Stan H.